OpenStudy (anonymous):

add sqrt2a+9sqrt8a^3

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

sqrt2a+9sqrt4*2a2*a =sqrt2a+18asqrt2a=(1+18a)sqrt2a

OpenStudy (anonymous):

can u draw that for me please?

OpenStudy (phi):

\[(1+18a)\sqrt{2a}\]

OpenStudy (anonymous):

thank you

jhonyy9 (jhonyy9):

yw